Jeannie and Ella work as sorters in a post office. If they work together,
they can finish the job in 24/7 hours. After working together for 2 hours,
Jeannie had a terrible headache and had to take a rest leaving Ella
for 2.5 hours more. How long will it take each one of them to do the work
alone?
HELP APPRECIATED... Thank you.

Answer by richwmiller(17219)

2/j+2/a+2.5/a=1
1/j+1/a=1/24/7
1/j+1/a=7/24
2/j+2/a+2.5/a=1
now we have
2/j+4.5/a=1,
1/j+1/a=7/24
a = 6, j = 8